OpenStudy (anonymous):

Let A be a nonsingular matrix. Show that det(A^-1) = 1/det(A)

OpenStudy (anonymous):

\[\det(A ^{-1})=\frac{ 1 }{ \det(A) }\] I'm assuming that it would have something to do with using the elementary matrices and how \[E_{k}...E_{1}A=I_{n}\] where \[E_{k}...E_{1}=A^{-1}\] but beyond that not sure how to set it up

OpenStudy (anonymous):

Determinant of a product of two matrices is the product of their determinants :D

OpenStudy (anonymous):

Yes. But, how would I show that this is true in general for all nonsingular matrices? I could set the determinant of each side equal to each other, in which case assuming that the claim is true, but still have no idea how I would show that it is actually true

OpenStudy (anonymous):

But \[\Large \det[AB]=\det[A]\det[B]\]?

OpenStudy (anonymous):

Yes. and I am told to PROVE that \[det(A^{-1}) = \frac{1} {det{A}}\] Wouldn't just saying that \[det(E_{k})...det(E_{1})det(A)=det(I_{n})\] be doing nothing more than assuming the initial statement as true and not prove anything?

OpenStudy (anonymous):

\[\Large A^{-1}A = I_n\]

OpenStudy (anonymous):

\[\Large \det[A^{-1}A]=\det[I_n]\]...

OpenStudy (anonymous):

\[\Large \det[I_n]=1\]\[\Large \det[A^{-1}A]= \det[A^{-1}]\det[A]\] stop me whenever ^.^

OpenStudy (anonymous):

okay. Sometimes I'm not quite sure how I've come this far in math...

OpenStudy (anonymous):

^.^